OpenStudy (anonymous):

write an equation of the line containing the given point and parallel to the given line express answer in y=mx+b (7,8) ;x+7y=4 y=

OpenStudy (saifoo.khan):

Parallel means same Slope.

OpenStudy (anonymous):

so what is the answer

OpenStudy (saifoo.khan):

now find b 8 = 7 (-1/7) + b 8 = -1 + b b = 9

OpenStudy (anonymous):

b=9 right

OpenStudy (saifoo.khan):

therefore: y = mx + b y = -1/7 x + 9

OpenStudy (saifoo.khan):

this is the ans.

OpenStudy (anonymous):

what about this one it is what is the function f(x)=2-there is two lines and an x in the middle like this /x/

OpenStudy (saifoo.khan):

this is known as MOD. in this, any value -ve or +ve... Always remains +ve.

OpenStudy (saifoo.khan):

| x |

OpenStudy (anonymous):

so it remains the same and how did u type that ?

OpenStudy (saifoo.khan):

No, it always remains +ve.

OpenStudy (anonymous):

o ok so what would it be and how do i work this problem out

OpenStudy (saifoo.khan):

to type that, there is a Pipe like thing above ENTER on your keyboard.

OpenStudy (saifoo.khan):

it is with the key of Back Slash.

OpenStudy (anonymous):

o ok

OpenStudy (saifoo.khan):

a guy drew a Graph there, he was correct about it.

OpenStudy (anonymous):

so how do we do this problem ?

OpenStudy (saifoo.khan):

well, ok look.

OpenStudy (saifoo.khan):

it is as simple as straight line graph. like the line will cut the y-axis @ 2. and its slope is -1. but as the line is going to touch x = 0, it will reflect back.

OpenStudy (saifoo.khan):

if u didnt undersand that, then i will use another way.

OpenStudy (anonymous):

ya use anothr way please

OpenStudy (anonymous):

so in this case it will be (2,-1) right

OpenStudy (saifoo.khan):

http://www.twiddla.com/569181

OpenStudy (saifoo.khan):

Join.

OpenStudy (anonymous):

can u help me with just two more please

OpenStudy (saifoo.khan):

Sure. Sure.

OpenStudy (saifoo.khan):

what was that MOD ques about? wht was the ques asking?

OpenStudy (anonymous):

a can we go back on that website and can u show me how to graph -1,-3 that goes with the equation y=-3-3x

OpenStudy (saifoo.khan):

OK!

OpenStudy (saifoo.khan):

Do u hav any more???

OpenStudy (saifoo.khan):

if not, thn i shall be leaving.